OpenStudy (luckycoins888):

fan + medal Find the simple interest. principal = $750; rate = 14%; time = 4.5 years A. $105 B. $420 C. $472.50 D. $1222.50

OpenStudy (anonymous):

I = prt Plug in the values I = 750*14%*4.5 \(\rightarrow\) I = 750*0.14*4.5

OpenStudy (anonymous):

So, multiplying starting from the left side of the equation, what is 750*0.14?

OpenStudy (luckycoins888):

105

OpenStudy (anonymous):

Correct, so what is 105*4.5?

OpenStudy (luckycoins888):

472.5